OpenStudy (anonymous):

I need help making a onto function that is not one to one so far i dont think a parabola will work

OpenStudy (anonymous):

@wilsondanielle

OpenStudy (wilsondanielle):

I lied i missed the second half of that question

OpenStudy (inkyvoyd):

I believe a parabola does work. Now it is up to you to show me why or why not.

OpenStudy (anonymous):

let me try something on wolfram

OpenStudy (wilsondanielle):

an onto function must use every y value. it doesn't need to pass the horizontal line test like a 1-1 function. Anything that passes the vertical line test would be an onto function.

OpenStudy (anonymous):

well a funtion like this f(x) = x^2 should work then

OpenStudy (anonymous):

What do you think @inkyvoyd @wilsondanielle https://www.desmos.com/calculator/lujt1dci93

OpenStudy (loser66):

How about y = x^3 -3x?

OpenStudy (anonymous):

that loos good, it is not 1 to 1

OpenStudy (loser66):

y= x(x^2 -3) has 3 roots x =0, \(x=\pm\sqrt3\) hence the graph looks like 

OpenStudy (loser66):

for every y, we have x but it is not one to one. Yeah !!

OpenStudy (anonymous):

https://www.desmos.com/calculator/v0vsfb0fxd

OpenStudy (wilsondanielle):

a parabola is fine, yes. It won't pass the horizontal line test (so is not 1-1) but it will pass the vertical line test and so will be an onto function.

OpenStudy (anonymous):

awesome!

OpenStudy (loser66):

Parabola is ok as long as you limit the range.

OpenStudy (anonymous):

next would be for me to find one that is neither onto or 1 to 1 lol f(x) = 0? think that will work?

OpenStudy (loser66):

Like \(f: R\rightarrow [0, \infty) \\x\mapsto x^2\)

OpenStudy (wilsondanielle):

that won't work. Pick a function with asymptotes - an onto function is continuous. Something like tan(x) would work.

OpenStudy (anonymous):

so something like f(x)= ceil(x/2)?

OpenStudy (anonymous):

oh i just saw that last part, are you say for neither onto or one to one or just onto?

OpenStudy (wilsondanielle):

tan(x) isn't onto, it would technically be 1-1 though as there are no duplicated y values for any given x value you would need a similar function to get neither onto nor 1-1. noncontinous = non onto multiple y values per x = non 1-1 find a function that satisfies both of those requirements.

OpenStudy (loser66):

Parabola works well for not one to one nor onto.

OpenStudy (anonymous):

i think i still know the function for a circle...

OpenStudy (wilsondanielle):

parabolas are onto. a circle is x^2+y^2

OpenStudy (loser66):

Like y = x^2 , not one to one for sure not onto also, since there exist y < 0 in the range but there is no x such that x^2 <0

OpenStudy (anonymous):

wouldn't that mean the last answer is not onto?

OpenStudy (wilsondanielle):

Parabolas are onto. The requirement for a function to be onto is not a y range from -infinity to infinity, but that every y value IN THE SET is used.

OpenStudy (anonymous):

so for the last answer the function y = x^2 would be good

OpenStudy (anonymous):

so isf(n)=⌈n/2⌉ onto, or one to one or both?

OpenStudy (wilsondanielle):

it is a linear function , so will be both

OpenStudy (anonymous):

so x = y^2 wont work either... https://www.desmos.com/calculator/hcx5srg6iy

OpenStudy (anonymous):

what about f(x) = 1/x-1 it is a discontinuous @wilsondanielle

OpenStudy (anonymous):

nope it wont do

OpenStudy (loser66):

the ceiling function is not one to one.

OpenStudy (anonymous):

What about f(x) =2/(x^2-x)

OpenStudy (loser66):

like ceil 3 =3, and ceil 2.6=3 also. Hence it is not one to one

OpenStudy (anonymous):

https://www.desmos.com/calculator/cxa3jfeppw

OpenStudy (anonymous):

what about that onto is the same as subjective right?

OpenStudy (anonymous):

@wilsondanielle

OpenStudy (loser66):

What do you want to find now?

OpenStudy (anonymous):

is the last graph i did good for not being one to one or onto?

OpenStudy (loser66):

Actually, It is good since some x's do not have a y. But I am not sure whether it is good to consider it is not onto or not. If they ask, which value of y you don't have preimage? the answer is "infinitive" which is so nonsense to me. :)

OpenStudy (anonymous):

lol

OpenStudy (loser66):

If I have y = | x| and I conclude that it is not one to one, since when \(x=\pm 1\), y = 1 It is not onto because if y = -1, I don't have any x such that y = |x|. Very simple and the proof is clear.

OpenStudy (anonymous):

ok im still lost, but i just got an idea

OpenStudy (loser66):

why lost?

OpenStudy (anonymous):

look at this: http://www.regentsprep.org/regents/math/algtrig/atp5/OntoFunctions.htm

OpenStudy (loser66):

so?

OpenStudy (anonymous):

according to that the parabola does not work if it is below 0 or a negetive

OpenStudy (loser66):

hey, look at "WHERE", that is the range. They limit the range to make it onto or not onto.

OpenStudy (anonymous):

i just checked it with wolfram it works for the problem

OpenStudy (loser66):

IN example 2, f : R to R, hence it is not onto

OpenStudy (anonymous):

even y = x^2 is not onto because it does not include every possble y value

OpenStudy (anonymous):

i see

OpenStudy (loser66):

In example 3 f : R to [-2, infinitive), hence it is onto.

OpenStudy (loser66):